OpenStudy (anonymous):

w varies jointly as x and y and inversely as the square of z. If w = 280 when x = 30, y = 12, and z = 3, find w when x = 20, y = 10 and z = 2.

OpenStudy (anonymous):

so w=kxy/(z^2) now in order to find w when x=20, y=10 and z=2, we must first find k or the proportionality constant... let's do that using the other information where w=280 if x=30, y=12 and z=3 manipulating this equation we'll have k=w(z^2)/xy k=280(3^2)/30*12 k=280*9/360 k=7 now that we know k, we can now compute for w when x = 20, y = 10 and z = 2: w=kxy/z^2 w=7*20*10/2^2 w=1400/4 w=350 <----the answer

OpenStudy (anonymous):

this is called a combined variation. Or simply a combination of the other kinds of variation or relationship

OpenStudy (anonymous):

^idk you but I think I love you.

OpenStudy (anonymous):

@rebeccaskell94 lol :))